# Preference Elicitation for Offline Reinforcement Learning

### v1

@@ Appendix C Environment Details
Figure 5: Star MDP illustrated in Figure 1. Transition probabilities are 0.9 for all solid arrows. Omitted actions or complementary transitions keep the state unchanged.
@@
We illustrate the transition dynamics underlying the Star MDP in Figure 5. Transition probabilities are 0.9 for all depicted solid arrows, and leave the state unchanged otherwise. Other actions also keep the state unchanged with probability 1. Episodes have length $H=3$ and start from $s_{0}$. Unless specified otherwise, the offline dataset $\mathcal{D}_{offline}$ consists of 40 trajectories which only cover states $(s_{0},s_{1},s_{3})$ and $(s_{3},s_{1},s_{2})$.
@@
The sepsis simulator [Oberst and Sontag, 2019] is a commonly used environment for medically-motivated RL work [Tang and Wiens, 2021]. We use the original authors’ publicly available code: https://github.com/clinicalml/gumbel-max-scm/tree/sim-v2/sepsisSimDiabetes (MIT license). The state space consists of five discrete observational variables (heart rate, blood pressure, oxygen concentration, glucose, diabetes status) and the action space consists of three different binary treatment options (antibiotic administration, vasopressor administration, mechanical ventilation). The probability that each treatment affects the value of each vital sign is determined by Oberst and Sontag [2019] to reflect patients’ physiology. The ground truth reward function is sparse and only assigns a positive reward of $+1$ to surviving patients and a negative reward of $-1$ if death occurs (3 or more abnormal vitals) during their stay. The offline trajectories dataset includes 10,000 episodes following an $\epsilon$-optimal policy with noise $\epsilon=0.1$ and the episode length is $H=20$.
Appendix D Additional Results
@@
Next, we plot performance as a function of preferences sampled for our two additional environments in Figure 8. We reach similar conclusion to those drawn from the Star MDP in Section 7: within the offline preference elicitation approaches, OPRL with uniform sampling is the least efficient, OPRL with uncertainty sampling performs better, and Sim-OPRL even better. The PbOP method naturally reaches a superior policy with fewer samples as it allows environment interaction and can thus improve its estimate of the transition model in parallel to learning the preference function.
(a) Gridworld.
(b) Sepsis simulation.
Figure 8: Empirical results on additional environments. Mean and 95% confidence interval over 20 experiments. Environment returns are normalised between 0 and 100. Only OPRL and Sim-OPRL are fully offline.
Appendix E Broader Impact

### v2

@@ Appendix D Environment Details
Figure 5: Star MDP. Transition probabilities are 0.9 for all solid arrows. Omitted actions or complementary transitions keep the state unchanged.
@@
We illustrate the transition dynamics underlying the Star MDP in Figure 5. Transition probabilities are 0.9 for all depicted solid arrows, and leave the state unchanged otherwise. Other actions also keep the state unchanged with probability 1. Episodes have length $H=3$ and start from $s_{0}$. Unless specified otherwise, the offline dataset $\mathcal{D}_{\textrm{offline}}$ consists of 40 trajectories which only cover states $(s_{0},s_{1},s_{3})$ and $(s_{3},s_{1},s_{2})$.
@@
Preferences collected over samples from the offline dataset learn slowly about the negative reward in the bottom state, as it is always included in the sampled trajectories. Instead, simulated rollouts can query a direct comparison between the optimal path and one that includes it. This example illustrates clearly why querying feedback over simulated rollouts achieves better environment returns than over samples from the offline buffer.
@@
MiniGrid-FourRooms.
We also conduct experiments on the minigrid-fourrooms-v0 D4RL dataset [Fu et al., 2020], ignoring all reward information in the offline dataset. In this environment, the agent must navigate a maze consisting of four interconnected rooms and reach a green goal square [Chevalier-Boisvert et al., 2023]. Agent and goal squares are randomly placed at the beginning of every episode.
HalfCheetah-Random.
The halfcheetah-random-v2 dataset is also part of the D4RL benchmark [Fu et al., 2020]. Our choice is motivated by Shin et al. [2022] who identify it as a particularly challenging offline preference-based reinforcement learning task. The dataset consists of 1 million transitions induced by a random policy in the MuJoCo environment Halfcheetah-v2, which rewards agents if they move forward. The observation space is 17-dimensional.
Table 4: Transition dynamics of the sepsis simulator environment [Oberst and Sontag, 2019]. Table adapted from [Tang and Wiens, 2021].
Step Variable Current New Change Variable Affected Effect
1 Antibiotics - on - Heart Rate high $\rightarrow$ normal w.p. 0.5
Systolic Blood Pressure high $\rightarrow$ normal w.p. 0.5
on off withdrawn Heart Rate normal $\rightarrow$ high w.p. 0.1
Systolic Blood Pressure normal $\rightarrow$ high w.p. 0.5
2 Ventilation - on - Oxygen Saturation low $\rightarrow$ normal w.p. 0.7
on off withdrawn Oxygen Saturation normal $\rightarrow$ low w.p. 0.1
3 Vasodilators - on - low $\rightarrow$ normal w.p. 0.7 (non-diabetic)
normal $\rightarrow$ high w.p. 0.7 (non-diabetic)
Systolic Blood Pressure low $\rightarrow$ normal w.p. 0.5 (diabetic)
low $\rightarrow$ high w.p. 0.4 (diabetic)
normal $\rightarrow$ high w.p. 0.9 (diabetic)
Blood Glucose very low, low $\rightarrow$ normal, normal $\rightarrow$ high,
high $\rightarrow$ very high w.p. 0.5 (diabetic)
on off withdrawn Systolic Blood Pressure normal $\rightarrow$ low w.p. 0.1 (non-diabetic)
high $\rightarrow$ normal w.p. 0.1 (non-diabetic)
normal $\rightarrow$ low w.p. 0.05 (diabetic)
high $\rightarrow$ normal w.p. 0.05 (diabetic)
4 Heart Rate fluctuate Vitals spontaneously fluctuate when not affected
5 Systolic Blood Pressure by treatment (either enabled or withdrawn),
6 Oxygen Saturation the level fluctuates ±1 w.p. 0.1, except:
7 Blood Glucose glucose fluctuates ±1 w.p. 0.3 (diabetic).
@@ 5 Systolic Blood Pressure by treatment (either enabled or withdrawn),
The sepsis simulator [Oberst and Sontag, 2019] is a commonly used environment for medically-motivated RL work [Tang and Wiens, 2021]. We use the original authors’ publicly available code: https://github.com/clinicalml/gumbel-max-scm/tree/sim-v2/sepsisSimDiabetes (MIT license). The state space consists of 1,440 discrete states based on different observational variables (heart rate, blood pressure, oxygen concentration, glucose, diabetes status). The action space consists of three binary treatment options (antibiotic administration, vasopressor administration, and mechanical ventilation). The complex transition dynamics of the environment determine how each treatment affects the value of each vital sign, and are summarized in Table 4; these were designed to reflect patients’ physiology [Oberst and Sontag, 2019]. The ground truth reward function is sparse and only assigns a positive reward of $+1$ to surviving patients and a negative reward of $-1$ if death occurs (3 or more abnormal vitals) during their stay. The offline trajectories dataset includes 10,000 episodes following an $\epsilon$-optimal policy with noise $\epsilon=0.1$ and the episode length is $H=20$.
Appendix E Additional Results
@@
Table 5: Sample complexity $N_{p}$ with Sim-OPRL as a function of dataset optimality, to reach a suboptimality gap of $\epsilon=20$ over normalized returns for the D4RL HalfCheetah environment. Mean and 95% confidence interval over 6 experiments.
Offline Dataset OPRL Uniform OPRL Uncertainty Sim-OPRL
HalfCheetah-Random 108 $\pm$ 9 71 $\pm$ 8 50 $\pm$ 10
HalfCheetah-Medium 63 $\pm$ 8 49 $\pm$ 6 36 $\pm$ 8
HalfCheetah-Medium-Expert 47 $\pm$ 8 45 $\pm$ 6 30 $\pm$ 8
Table 6: Policy performance for a fixed preference budget ($N_{p}=30$) for different D4RL HalfCheetah datasets. Normalized environment returns, mean and 95% confidence interval over 6 experiments.
Offline Dataset OPRL Uniform OPRL Uncertainty Sim-OPRL
HalfCheetah-Random 17 $\pm$ 7 47 $\pm$ 10 49 $\pm$ 9
Halfcheetah-Medium 43 $\pm$ 6 63 $\pm$ 6 69 $\pm$ 7
Halfcheetah-Medium-Expert 57 $\pm$ 6 72 $\pm$ 6 80 $\pm$ 6
To complement the analysis with a more complex environment, we also ran different preference elicitation algorithms on D4RL HalfCheetah datasets of varying optimality. In Table 6, we report the number of preference samples needed to achieve a target suboptimality with these different datasets. We reach the same conclusion as above: fewer preferences are needed as the dataset becomes more optimal.
In Table 6, we report policy performance for a fixed preference budget. With a fixed preference budget, policies learned from suboptimal observational datasets achieve lower returns, due to the worse quality of the transition model. Shin et al. [2022] reach similar conclusions when evaluating policy performance under different OPRL methods (in their Table 2).
Appendix F Broader Impact